A. Areas under the Normal Curve: Find the area of the following. Then, illustrate using the

normal curve.

1. 𝑧 = 0.34

2. 𝑧 = 2.12

3. 𝑧 = βˆ’1.35

4. 𝑧 = βˆ’0.27

5. 𝑧 = 1.07

6. At least 𝑧 = βˆ’0.47

7. Between 𝑧 = 0.76 π‘Žπ‘›π‘‘ 𝑧=2.34

8. Greater than 𝑧 = 0.78

9. Less than 𝑧 = βˆ’0.67

10. Between 𝑧 =βˆ’1.52 π‘Žπ‘›π‘‘ 𝑧=0.97


1
Expert's answer
2022-04-27T00:55:26-0400

1. 0.6331



2.0.9830



3. 0.0885



4.0.3936



5.0.8577



6.1-0.3192=0.6808



7.0.9904-0.7764=0.214



8.1-0.7823=0.2177



9.0.2514



10.0.8340-0.0643=0.7697